Welcome >  Stations >  Zimmerwald, Switzerland (7810) >  Station Logs

Station History Log - Zimmerwald, Switzerland (7810)

SOD Year Day
of Year
Time
of Day
Data Impact System Description
78106801 2021 341 08:00 0 06 RECEIVER Discriminator-Eventtimer: For the event timer start1, connected two NIM outputs of the discriminator in parallel. This fixed the high calibration RMS due to changed earthing and grounding since install of STATION UPS (doy= 167).
78106801 2021 293 08:00 0 05.09 LASER IR-attenuator re-adjusted. Energies(RegenOut,IR-TP,GROut) 3.8 mJ, 21 mJ, 8.2 mJ. End of manufacturer service.
78106801 2021 292 14:00 1 05.09 LASER Regeneration amplifier: Medox-Triode replaced because the manufacturer technician is at site. Old triode still good. With new triode 1.step is at -4kV (entire HV is at ~-9.5kV) adjusted for max. pulse energy in the green.
78106801 2021 292 12:00 0 05.09 LASER Mirror M8 replaced because of a small damage.
78106801 2021 292 10:00 0 05.09 LASER Medox-Delay2(D13) 270.23250us increased pulse energy by reducing post pulses and select previous regen pulse.
78106801 2021 291 14:00 0 05.09 LASER Medox high voltage adjusted in order to compensate for better voltage symmetry at pockels cell around switching point. Reduced post-pulses and increased pulse energy in the green.
78106801 2021 256 09:00 1 05.09 LASER Compensated start diode time walk/Calibration constant C3 changed by +1ns in order to move cal histogram to 0.
78106801 2021 256 09:00 1 05.09 LASER Compensation of lower gain: Startdiode: Discriminator: Threshold reduced from 0.600 V to 0.300 V and at PowerDivider: 50 Ohm terminator removed.
78106801 2021 229 13:00 0 05 LASER Chiller: Particle filter exchanged (after too low flow).
78106801 2021 203 18:00 0 99 COOLING operator room: Installed new outdoor device. Deinstalled temporary replacement air condition.
78106801 2021 173 10:00 0 06 RECEIVER Discriminator-Eventtimer: For the event timer stop1, connected two NIM outputs of the discriminator in parallel.
78106801 2021 167 12:00 2 99 STATION UPS: Installed one new for the entire station. Removed UPS SLR1/2. This changed slightly earthing and grounding.
78106801 2021 152 11:45 0 05.09 LASER Medox-Delay2(D13) 270.24750us increased main pulse by reducing post pulse and select next regen pulse.
78106801 2021 145 10:00 0 06 RECEIVER Fabry-Perot: micro meter screws position GR:66.6(new scale) adjusted using Coude path retro reflector simulation.
78106801 2021 124 10:00 0 06 RECEIVER C-SPAD: For cross hair adjustment in front of the detector, slightly changed the vertical beam alignment of receive mirror RM6.
78106801 2021 116 10:00 0 08.02 CAL: End of range bias. There was not yet a real fix concerning high RMS. So, the problem might still remain.
78106801 2021 113 10:00 2 08.02 CAL: Beginning of range bias due to a coaxial tee and cable workaround at the event timer stop input to get stops again.
78106801 2021 096 15:00 0 08.02 CAL: End of high outliers? Temperature gradient around the detector box reduced by rearranging the replacement air condition.
78106801 2021 050 00:00 2 08.02 CAL: Beginning of high outliers. An air condition was gone. A replacement air condition was setup.
78106801 2020 344 13:00 0 05 LASER Chiller: DI-cartridge exchanged (after too high conductivity).
78106801 2020 325 13:00 0 06 RECEIVER Fabry-Perot: micro meter screws position GR:20.5,IR:71.8, IR adjusted using Coude path retro reflector simulation.
78106801 2020 294 10:00 0 05.09 LASER Medox-Delay2(D13) 270.23350us undone last change.
78106801 2020 253 10:00 0 05.09 LASER Medox-Delay2(D13) 270.24750us increased main pulse by reducing post pulse and select next regen pulse.
78106801 2020 212 18:00 0 4.01 TELESCOPE Elevation Encoder: Cleaned scanning head due to high pollution.
78106801 2020 196 07:00 0 09 TIME+FREQ Set maser frequency synthesiser value to 869.314638.
78106801 2020 163 10:00 0 05 LASER Infrared: Removed IR foam/absorber located after beam splitter/polarizer P10 and before P11.
78106801 2020 162 13:00 0 05 LASER Chiller: Particle filter exchanged (after too low flow).
78106801 2020 093 13:00 0 05 LASER beam direction alignment within telescope socket
78106801 2020 007 07:45 0 09 TIME+FREQ Set maser frequency synthesiser value to 869.314505 and 1PPS(Z) to GPS-30 us.
78106801 2019 182 13:00 1 05.09 LASER Compensated start diode time walk/Calibration constant C3 changed by -20ps/6mm in order to move cal histogram to 0.
78106801 2019 169 18:00 1 05.09 LASER Regeneration amplifier: Medox-Triode replaced due to too high pre-pulses
78106801 2019 092 15:50 0 09 TIME+FREQ Set maser frequency synthesiser value to 869.314352 and 1PPS(Z) to GPS-30 us.
78106801 2019 058 18:00 0 09 TIME Local time scale: local 1PPS offsets are now measured using a Trimble GPS receiver and a Standfort counter.
78106801 2018 299 13:05 0 09 TIME+FREQ Set maser frequency synthesiser value to 869.314150 and 1PPS(Z) to GPS-30 us.
78106801 2018 226 08:00 0 05 LASER Chiller: DI-cartridge and particle filter exchanged (after too high conductivity).
78106801 2018 058 10:55 0 12.03 METEO Humidity Sensor: After 100 % humidity, showed a memory effect and needed too long time to decrease to the real value. Replaced sensor by another model.
78106801 2017 325 17:11 0 09 TIME Local time scale: Between 13:11 and 17:11 generated time bias.
78106801 2017 304 07:49 0 09 TIME Local time scale: Locked to GPS again and, thus, time bias at usual GPS offsets again.
78106801 2017 296 13:11 0 09 TIME Local time scale: lost due to power loss of the garage USV24V and the connected devices. Generated huge time bias.
78106801 2017 166 07:45 0 09 TIME+FREQ Set maser frequency synthesiser value to 869.314070 and 1PPS(Z) to GPS-30 us.
78106801 2017 061 18:00 0 06 RECEIVER Fabry-Perot: micro meter screws position 520, using Coude path retro reflector simulation.
78106801 2017 052 18:00 1 05.09 LASER Compensated start diode time walk
78106801 2017 038 18:00 0 05.09 LASER Medox-Delay2(D13) 270.23350us using new double pass amplifier
78106801 2017 038 18:00 1 05.08 LASER Double pass amplifier: DPCH 256 replaced by new DPCH 236, aligned, regen re-adjusted
78106801 2017 019 18:00 0 05 LASER Chiller: DI-cartridge and particle filter exchanged (after too high conductivity).
78106801 2016 330 08:30 0 09 TIME+FREQ Set maser frequency synthesiser value to 869.313893 and 1PPS(Z) to GPS-30 us.
78106801 2016 316 18:00 0 05.08 LASER Double pass amplifier: DPCH 256 removed (to be replaced be a new DPCH 236 in the near future).
78106801 2016 295 18:00 0 05.08 LASER Medox Driver: High voltage power supply defect
78106801 2016 258 10:30 0 09 TIME+FREQ Set maser frequency synthesiser value to 869.313863 and 1PPS(Z) to GPS-30 us.
78106801 2016 196 18:00 0 05.08 LASER Double pass amplifier: Second compensation of lower gain: Startdiode: Discriminator: Threshold reduced from 0.600 V to 0.300 V.
78106801 2016 180 18:00 0 05.08 LASER Double pass amplifier: Second lower gain.
78106801 2016 123 18:00 0 05.08 LASER Double pass amplifier: First compensation of lower gain: Startdiode: PowerDivider: 50 Ohm terminator removal.
78106801 2016 118 18:00 0 05.08 LASER Double pass amplifier: First lower gain.
78106801 2016 099 18:00 0 05.08 LASER Regeneration amplifier: For optimisation, M5 and pockels cell adjustment at 25 degree water temperature
78106801 2016 099 18:00 0 05 LASER Chiller: For regen stability, increased operating water temperature from 20 to 22 degrees. SLR operation continues
78106801 2016 078 18:00 0 05 LASER Chiller: For energy maximisation, reduced operating water temperature from 25 to 20 degrees
78106801 2016 077 18:00 1 05.08 LASER Double pass amplifier: DPCH 256 re-installed after repair
78106801 2015 338 18:00 0 05.08 LASER Double pass amplifier: DPCH 256 removed for repair.
78106801 2015 294 18:00 0 05.08 LASER Regeneration amplifier: For stabilisation, optimisation of the Regen and seeding
78106801 2015 258 18:00 0 05.08 LASER Regeneration amplifier: Measured output energy: 3.4 mJ
78106801 2015 258 18:00 0 05 LASER Chiller: DI-cartridge and particle filter exchanged (after too high conductivity), now 0.2uS instead of 12uS. Water connectors at laser box exchanged.
78106801 2015 134 18:00 0 06 RECEIVER Fabry-Perot: micro meter screws position 497, new optics alignment compensated, raw adjustment
78106801 2015 098 18:00 1 05.08 LASER Regeneration amplifier: DPCH 236 re-installed after repair
78106801 2015 098 18:00 0 05.08 LASER Medox-Delay2(D13) 270.23400us using repaired regeneration amplifier
78106801 2015 059 18:00 0 99 ZIMLAT-PC PFI=1 (derotates now again); was deactivated due to maintainance
78106801 2015 021 18:00 0 4.01 TELESCOPE Adjustment: Adjustment support removed from elevation axis. We had CCD observations through paper at the SI1100 camera
78106801 2015 013 18:00 0 4.01 TELESCOPE Adjustment: Integrated adjustment support at the elevation axis
78106801 2015 011 18:00 0 4.01 TELESCOPE Adjustment: Adaptation of the adjustment laser beam from the optical bench towards the optical axis of the telescope
78106801 2015 007 18:00 0 99 CAMERA Shutter of the SI1100: evaluated open/close delays at the freezer: min/typ/max=50ms/55ms/65ms
78106801 2014 306 18:00 0 05.08 LASER Regeneration amplifier: Water leakage, shut down SLR-operation
78106801 2014 290 18:00 0 99 STATION Cooling: Water filled-up
78106801 2014 287 18:00 0 05 LASER Chiller: Manufacturer replaced defect hot gas/magnete valve; DI-cartridge and particle filter exchanged too
78106801 2014 262 18:00 0 05 LASER Thick laser head cable moved/kneaded in order to get current pulses without pulse-end ripples again
78106801 2014 259 18:00 0 99 STATION Cooling: Water filled-up
78106801 2014 259 18:00 0 10.01 DATAEVAL normpt.f: Fixed the normal point bin rms by correcting statistics
78106801 2014 252 18:00 0 99 DOME Engines and reduction gear exchanged
78106801 2014 204 18:00 0 05.08 LASER Medox-Delay2(D13) 270.24800us using new double pass amplifier
78106801 2014 203 00:00 1 05.07 LASER Double pass amplifier: DPCH 236 replaced by DPCH 256, Medox-Triode replaced
78106801 2014 154 18:00 0 05 LASER Chiller: DI-cartridge and particle filter exchanged (after too high conductivity), now 0.4uS..0.9uS instead of 13uS
78106801 2014 028 18:00 0 05 LASER Chiller: Conductivity feedback loop activated, conductivity-set value range set to standard (4uS..7uS)
78106801 2014 015 18:00 0 05.07 LASER Medox-Delay2(D13) 270.23300us read out only
78106801 2014 015 18:00 0 05.07 LASER Medox-Delay1(D12) 270.00400us read out only
78106801 2013 349 18:00 0 05 LASER Chiller: De-ionisation cartridge exchanged after too high conductivity, now 0,4uS instead of 8?uS, Reason: usually exchange after two years, but a lot of changes in the water loop: manometer integrated and removed at different locations, 1. flow sensor replaced, 2. flow sensor integrated, laser heads unplugged and re-plugged
78106801 2013 315 18:00 0 05 LASER Chiller: Particle filter: screwed more tightly in order to loose no more water
78106801 2013 289 18:00 0 05 LASER Chiller: Flow sensor replaced. Posthumous: This was the reason for the chiller flow errors
78106801 2013 261 18:00 0 05 LASER Chiller: De-ionisation cartridge exchanged after too high conductivity, now 0,4uS instead of 4uS
78106801 2013 200 18:00 0 05 LASER Chiller: Water filled-up to max after emptying completely
78106801 2013 200 18:00 0 05 LASER Chiller: Water completely removed und hoses of the secondary loop flushed using destillated water
78106801 2013 200 18:00 0 05 LASER Chiller: Particle filter exchanged, flow before/after exchange <5,0 (error flow 1) / 8,3 l/min
78106801 2013 196 18:00 0 05 LASER Chiller: Water filled-up to max due to conductivity warning alarm
78106801 2013 185 18:00 0 05.07 LASER Medox-Delay2(D13) 270.23305us using repaired double pass amplifier
78106801 2013 185 18:00 0 05.07 LASER Double pass amplifier (Replacement head): DPCH236 installed after repair of a diode
78106801 2013 185 18:00 0 05 LASER Chiller: Water filled-up to max due to integration of the repaired laser head two replacements
78106801 2013 161 18:00 0 05 LASER Chiller: Particle filter full with particles (determined after laser head two replacement has now malfunction)
78106801 2013 138 18:00 0 05 LASER Chiller: Water filled-up to max (after alarm)
78106801 2013 122 18:00 0 06 RECEIVER Fabry-Perot: micro meter screws position 550, Fabry-Perrot positioning support integrated, vernier adjustment
78106801 2013 100 18:00 0 05 LASER Chiller: Water filled-up to max due Particle filter exchange
78106801 2013 100 18:00 0 05 LASER Chiller: Particle filter exchanged, flow before/after exchange <7,0 (error flow 1) / 8,3 l/min
78106801 2013 077 18:00 0 05 LASER Chiller: Water filled-up to max (after alarm) again after three weeks already
78106801 2013 056 18:00 0 05 LASER Chiller: Water filled-up to max (after alarm)
78106801 2013 045 18:00 0 05.07 LASER Medox-Delay2(D13) 270.2480 us Pre-/Post-Pulse Correction
78106801 2013 045 00:00 1 05.06 LASER Double path amplifier: DPCH 256 replaced by DPCH 236
78106801 2013 044 18:00 0 05.06 LASER Medox-Delay2(D13) 270.24780us read out only
78106801 2013 016 00:00 1 05.05 LASER SHG crystal replaced
78106801 2012 346 18:00 0 05 LASER Chiller: Particle filter exchanged (Device manufacturer repaired device and exchanged the filter)
78106801 2012 345 18:00 0 05.05 LASER Double path amplifier: Water leakage
78106801 2012 316 18:00 0 99 STATION Cooling: Water filled-up
78106801 2012 300 18:00 0 05 LASER Chiller: Particle filter exchanged
78106801 2012 027 00:00 1 05.04 RECEIVER Event Timer: ET32 #3203 replaced by ET32 #3207
78106801 2011 181 18:00 0 05.04 LASER Medox-Delay2(D13) 270.2845 us read from documentation
78106801 2010 264 18:00 0 06 RECEIVER Fabry-Perot: micro meter screws position 633
78106801 2010 243 18:00 0 05 LASER Chiller: Particle filter exchanged, flow before/after exchange 7,5 / 8,3 l/min
78106801 2008 197 18:00 0 06 RECEIVER Fabry-Perot: micro meter screws position 700, SPAD re-adjusted
78106801 2008 085 00:00 3 05 STATION New baseline configuration
78106801 2008 049 18:00 0 05.03 LASER MEDOX replaced. Effect: Lower single shot rms: (50 ps for 423 nm SPAD calibration: 80 ps for 423 nm SPAD Lageos)
78106801 2006 172 00:00 2 08.02 CAL Pre- and after-pass calibration to terrestrial target
78106801 2006 082 00:00 3 05 RECEIVER 423 nm: Hamamatsu PM + A032ET / 846 nm: PM + A032ET
78106801 2006 034 00:00 3 05 RECEIVER 423 nm: Hamamatsu PM + A032ET / 846 nm: PM + SR620 timer
78106801 2003 248 18:00 0 05.02 LASER New etalon installed on 9 March 2001
78106801 2001 068 18:00 0 05.01 LASER Pulse width was not according to specs: (verified with streak camera end of 2000)
78106801 2001 067 00:00 2 05 LASER Laser pulse width reduced to 100 ps FWHM
78106801 1999 236 00:00 3 10.01 DATAEVAL Normalpoint bin shifted by half an interval to conform
78106801 1997 355 00:00 3 05 RECEIVER Detector Hamamatsu PM replaced. New: H6533
78106801 1997 001 00:00 3 05 STATION Baseline configuration

Find more topics on the central web site of the Technical University of Munich: www.tum.de